OpenStudy (anonymous):

Ok I give up. I just need the answers please to get by this section. (-2, -1) & (-9, -4) in slope intercept form.

OpenStudy (anonymous):

m = (y2 - y1)/(x2 - x1) so, (-4 +1)/(-9 + 2) you get -3/-7 = 3/7 so you have ur slope intercept formula y = mx + b you got ur m = 3/7 y = (3/7)x + b now choose any of the 2 points (doesnt matter which), ill choose (-2, -1) plug and chug -1 = (3/7)(-2) + b solve for b -1/7 = b so your answer is y = (3/7)x -1/7